# Deep Learning Framework for History Matching CO2 Storage with 4D Seismic and Monitoring Well Data

### v1

@@ 5.2.2 Monitoring data surrogate evaluation
Figure 10: Comparison of time evolution between monitoring well surrogate saturation predictions (surr) and reference simulation results (sim) for three realizations. Each column corresponds to one of the 16 time steps considered.
@@
$\delta_{s}=\frac{1}{n}\sum_{i=1}^{n}\lvert S_{i}-\hat{S}_{i}\rvert,$ (12)
where $S_{i}$ and ${\hat{S}}_{i}$ denote the reference and surrogate model prediction, respectively, and $n$ is the total number of contributions to $S_{i}$ and ${\hat{S}}_{i}$. For the seismic surrogate we have $n=6\times 42\times 42\times 7=74,088$, and for the monitoring well surrogate model $n=16\times 35=560$. These errors are computed for each test case.
The MAEs for both the interpreted seismic surrogate and the monitoring well surrogate, over the 500 test-case realizations, are shown as box plots in Figure 11. In the box plots, the top and bottom of the box represent the 75th and 25th percentile (P75 and P25) errors and the lines outside the boxes represent the 90th and 10th percentile errors. The red lines inside the boxes correspond to the median error. These value are 0.00384 and 0.0155 for interpreted seismic data and monitoring well data, respectively. These low error values suggest that the surrogate models are sufficiently accurate for use in history matching.
Figure 11: Box plots of mean absolute saturation errors for interpreted seismic surrogate (Seismic) and monitoring well surrogate (Monitor) over 500 test-case models.
@@ 6.1 Setup, errors and likelihood function
The true (synthetic) geomodel, used to provide observed data, is a realization characterized by the following metaparameter values: $\mu_{\log k}=3.0$, $\sigma_{\log k}=2.3$, $\log_{10}a_{r}=-1.8$ ($a_{r}=0.016$), $d=0.03$ and $e=0.055$. The true-model response for CO2 saturation, along $y$-$z$ and $x$-$z$ cross sections through the injection well at a time of 1 year, is presented in Figure 12(a) and (b). The corresponding filtered seismic interpretations are displayed in Figure 12(c) and (d). The filtered response loses some of the detail, but the heterogeneity-driven channeling features still appear.
@@
Figure 12: True-model saturation fields at 1 year. Top row shows high-fidelity results for cross sections through the injection well, and bottom row shows corresponding filtered (interpreted seismic) results.
Noise, intended to represent measurement error, model resolution error, and surrogate approximation error, is added to the true data. The observed data vector is thus expressed as:
$\mathbf{d}_{\text{obs}}=\mathbf{d}_{\text{true}}+\boldsymbol{\epsilon},$ (13)
where $\boldsymbol{\epsilon}$ denotes all sources of noise. Here we take $\boldsymbol{\epsilon}$ to be uncorrelated and unbiased, i.e., of zero mean and diagonal covariance matrix. The measurement error and model resolution error are assigned directly in this work. For the monitoring well data, the standard deviation of the combined measurement and model resolution errors is set to be 5% of the maximum of the collected monitoring data over all available time steps. In our case this is 0.0162, which is close to the value (0.02) used by Sun and Durlofsky (2019). For the interpreted seismic data, the standard deviation of the error is set to be 10% of the maximum of the observed data over all available time steps. This is similar to the approach used by Gervais and Roggero (2010), who set the standard deviation of error to be 10% of the seismic data mean.
The surrogate-model approximation error ($\boldsymbol{\epsilon}_{\text{surr}}$) is calculated during the test-set evaluation. Specifically,
@@
where $\mathbf{d}_{\text{pred}}$ and $\mathbf{d}_{\text{true}}$ denote data predicted by the surrogate and true models, respectively. In our case, the mean and standard deviation of error in the monitoring data surrogate are 0.00124 and 0.0370, respectively. For the seismic surrogate, these values are -0.0001 and 0.0133, respectively. Thus we see that the average errors are near zero, consistent with the assumption of unbiased errors. The total standard deviation ($\sigma_{\text{total}}$), which combines the standard deviation of $\boldsymbol{\epsilon}_{\text{surr}}$ ($\sigma_{\text{surr}}$) and the measurement and model resolution errors (we denote the standard deviation of these two errors as $\sigma_{\text{d}}$), can then be obtained via $\sigma_{\text{total}}=\left(\sigma_{\text{surr}}^{2}+\sigma_{\text{d}}^{2}\right)^{1/2}$. The noise for the interpreted seismic data is added to the ‘true’ high-fidelity simulation results before the filtering procedure, consistent with the approach used by Bukshtynov et al. (2015). Note that other treatments for error can be readily incorporated into our framework. An approach that might be considered is one in which error varies as a function of saturation value, as could be the case if seismic data are more sensitive in some saturation ranges than others.
@@
We will apply the hierarchical MCMC method introduced in Section 4 for cases involving both monitoring data and seismic data, and for cases involving only monitoring data. This will allow us to assess the impact of seismic data on uncertainty reduction. In the former case, the likelihood appearing in Eq. 10 and Eq. 11 is given by
@@
$\begin{split}p(\mathbf{d}_{\text{obs}}\mid\boldsymbol{\xi},\mathbf{h})=c\exp\left(\frac{1}{T}\left(-\frac{1}{2}(\mathbf{d}_{\text{obs}}^{\text{mon}}-f_{\text{mon}}(\boldsymbol{\xi},\mathbf{h}))^{T}C_{\text{mon}}^{-1}(\mathbf{d}_{\text{obs}}^{\text{mon}}-f_{\text{mon}}(\boldsymbol{\xi},\mathbf{h}))\right.\right.\\ \left.\left.-\frac{1}{2}(\mathbf{d}_{\text{obs}}^{\text{seis}}-f_{\text{seis}}(\boldsymbol{\xi},\mathbf{h}))^{T}C_{\text{seis}}^{-1}(\mathbf{d}_{\text{obs}}^{\text{seis}}-f_{\text{seis}}(\boldsymbol{\xi},\mathbf{h}))\right)\right),\end{split}$ (15)
@@
where $c$ is a normalization constant, $\mathbf{d}_{\text{obs}}^{\text{mon}}$ and $\mathbf{d}_{\text{obs}}^{\text{seis}}$ denote observed monitoring well data and interpreted seismic data, respectively, and $C_{\text{mon}}$ and $C_{\text{seis}}$ are the (diagonal) total error covariance matrices for monitoring well and seismic data, respectively. The quantity $T$ ($T>0$), referred to as simulated temperature, is a parameter that controls the acceptance rate of proposed samples (Li, 2012). If only the monitoring data are used for history matching, the seismic data portion of the likelihood does not appear, resulting in
@@
$\begin{split}p(\mathbf{d}_{\text{obs}}\mid\boldsymbol{\xi},\mathbf{h})=c\exp\left(\frac{1}{T}\left(-\frac{1}{2}(\mathbf{d}_{\text{obs}}^{\text{mon}}-f_{\text{mon}}(\boldsymbol{\xi},\mathbf{h}))^{T}C_{\text{mon}}^{-1}(\mathbf{d}_{\text{obs}}^{\text{mon}}-f_{\text{mon}}(\boldsymbol{\xi},\mathbf{h}))\right)\right).\end{split}$ (16)
@@
Additional parameter specifications are as follows. We set $\beta=0.05$, where $\beta$ (appearing in Eq. 9) scales the magnitude of sample updates. The standard deviation of the Gaussian proposal distribution for the metaparameters is set to be 0.05 of the parameter range. We specify $T=1$ when only monitoring data are used, and $T=5$ when both data types are used. These settings result in MCMC acceptance rates in the appropriate range, i.e., 10% to 40% (Gelman et al., 1996; Han et al., 2024). In the results below, hierarchical MCMC using only monitoring well data requires 47,569 function evaluations (18,749 models accepted). Results using both data types require 55,748 function evaluations (16,249 models accepted).
@@ 6.2 History matching results
Posterior results for the metaparameters, obtained from the last 10,000 accepted samples, for cases both with and without seismic data, are presented in Figures 13 and 14. The gray region in each figure represents the prior distribution of the metaparameter and the red dashed line indicates the true value of the metaparameter. The posterior histograms appear in blue for results in which only monitoring well data are used, and in orange for results where both monitoring well data and interpreted seismic data are used. A reasonable degree of uncertainty reduction is achieved for $\mu_{\log k}$, $\sigma_{\log k}$ and $\log_{10}a_{r}$ when only monitoring well data are used (Figure 13(a), (c) and (e)). In addition, it is evident that the mode of the posterior histogram is near the true value for these quantities.
Very little uncertainty reduction is observed for parameters $d$ and $e$ in this case (Figure 14(a) and (c)), indicating that the monitoring well data are not informative in terms of the relationship between $\phi$ and $\log k$. This is consistent with the observations of Han et al. (2024). Uncertainty reduction is achieved, however, in the porosity field itself. This is demonstrated in Figure 14(e), where we show the prior and posterior distributions for the mean of the porosity, $\mu_{\phi}$. The results in Figure 14(e) are determined by computing, in turn, the average porosity for 3000 randomly sampled prior and posterior realizations.
Considerably more uncertainty reduction in the key metaparameters is achieved when both monitoring well data and seismic data are used. Specifically, we observe narrower distributions, which include the true value, for $\mu_{\log k}$, $\sigma_{\log k}$ and $\log_{10}a_{r}$ (Figure 13(b), (d), (f)). There is also more uncertainty reduction in $d$ and $e$ in this case (Figure 14(b) and (d)), though significant posterior uncertainty in these quantities remains. As is apparent in Figure 14(f), the use of both data types provides substantial uncertainty reduction in mean porosity itself. Taken in total, the results in Figures 13 and 14 clearly show the additional (significant) uncertainty reduction achieved when interpreted seismic data are incorporated into the data assimilation procedure.
@@
(b) $\mu_{\log k}$ with seismic
(c) $\sigma_{\log k}$ without seismic
(d) $\sigma_{\log k}$ with seismic
(e) $\log_{10}a_{r}$ without seismic
(f) $\log_{10}a_{r}$ with seismic
Figure 13: History matching results for metaparameters $\mu_{\log k}$, $\sigma_{\log k}$ and $\log_{10}a_{r}$ with monitoring well data only (left column, blue histograms) and with both monitoring well data and interpreted seismic data (right column, orange histograms). Priors shown in gray and true values indicated by the vertical red line.
@@
(b) $d$ with seismic
(c) $e$ without seismic
(d) $e$ with seismic
(e) $\mu_{\phi}$ without seismic
(f) $\mu_{\phi}$ with seismic
Figure 14: History matching results for metaparameters $d$ and $e$ and average porosity $\mu_{\phi}$ with monitoring well data only (left column, blue histograms) and with both monitoring well data and interpreted seismic data (right column, orange histograms). Priors shown in gray and true values indicated by the vertical red line.
We next present saturation fields (CO2 plumes) for prior and posterior geomodels. In order to identify ‘representative’ interpreted seismic saturation fields, and thus enable meaningful comparisons, we proceed as follows. We randomly select 1000 prior geomodel realizations from the training set, 1000 realizations from the geomodels accepted during history matching with only monitoring data, and 1000 realizations from the geomodels accepted during history matching with both data types. Interpreted seismic data are then generated for each of these 3000 realizations using the seismic surrogate model. Then, for each set of models (prior, history matching with monitoring data, history matching with both data types), we apply k-means clustering to construct four clusters. The medoid from each cluster is then viewed as a representative interpreted seismic saturation field.
The resulting 12 saturation fields are shown in Figure 15 ($y$-$z$ cross sections through the injection well) and in Figure 16 ($x$-$z$ cross sections through the injection well). These results are all at a time of 1 year. The true results for this case are shown in Figure 12(c) and (d). The plumes in the prior realizations are fairly regular, varying from near-cylindrical to cone-shaped. The plumes for the posterior models all differ considerably from any of the prior plumes, and all resemble the true result to some degree. Specifically, in both figures, the posterior plumes all display CO2 channeling toward the bottom of the model. Posterior results in which both data types are used are closer to the true saturation fields. For example, in Figure 15(i)-(l) we see three somewhat distinct channels to the left of the injector along with CO2 moving further to the left than the right. These behaviors, consistent with Figure 12(c), are not consistently observed in Figure 15(e)-(h). In addition, the extent of CO2 channeling toward the bottom of the model is better captured in Figure 16(i)-(l) than in Figure 16(e)-(h) (compare to Figure 12(d)). These results reiterate the advantages of using both data types for history matching.
@@
(i) Post. 1 (with seis)
(j) Post. 2 (with seis)
(k) Post. 3 (with seis)
(l) Post. 4 (with seis)
Figure 15: Representative interpreted seismic saturation fields in $y$-$z$ cross sections through the injector at 1 year for prior realizations (first row), posterior realizations obtained using only monitoring well data (second row), and posterior realizations obtained using both monitoring well and seismic data (third row). True result shown in Figure 12(c).
@@
(i) Post. 1 (with seis)
(j) Post. 2 (with seis)
(k) Post. 3 (with seis)
(l) Post. 4 (with seis)
Figure 16: Representative interpreted seismic saturation fields in $x$-$z$ cross sections through the injector at 1 year for prior realizations (first row), posterior realizations obtained using only monitoring well data (second row), and posterior realizations obtained using both monitoring well and seismic data (third row). True result shown in Figure 12(d).

### v2

@@ 5.2.2 Monitoring well data surrogate evaluation
Figure 11: Comparison of time evolution between monitoring well surrogate saturation predictions (surr) and reference simulation results (sim) for three realizations. Each column (in each subfigure) corresponds to one of the 16 time steps considered.
@@
$\delta_{s}=\frac{1}{n}\sum_{i=1}^{n}\lvert S_{i}-\hat{S}_{i}\rvert,\ \ {\rm for}\ S_{i}>\varepsilon\ {\rm or}\ {\hat{S}}_{i}>\varepsilon,$ (12)
where $S_{i}$ and ${\hat{S}}_{i}$ denote the reference and surrogate model prediction, respectively. Here the contribution to $\delta_{s}$ is calculated only for grid cells where either simulated or predicted saturation is higher than a threshold $\varepsilon$. Consistent with this, $n$ is the total number of cells where $S_{i}>\varepsilon$ or ${\hat{S}}_{i}>\varepsilon$. Thus, the value of $n$ will be different for different realizations. A value for $\delta_{s}$ is computed for each test case for both surrogate models.
The MAEs for both the interpreted seismic surrogate and the monitoring well surrogate, over the 500 test-case realizations, are shown as box plots in Figure 12. The various boxes for each case correspond to different thresholds ($\varepsilon=0$, 0.01, 0.02, 0.05). In the box plots, the top and bottom of the box represent the 75th and 25th percentile (P75 and P25) errors and the lines above and below the boxes represent the 90th and 10th percentile errors. The orange lines inside the boxes correspond to the median error. The median MAEs range from 0.006 to 0.023 for interpreted seismic data, and from 0.019 to 0.028 for monitoring well data. Error is lowest for $\varepsilon=0$ because many cells with near-zero saturation (and thus near-zero error) contribute to the MAE calculation. The errors remain fairly consistent and relatively low for different nonzero threshold values. As we will see, this level of accuracy is more than sufficient to enable the use of the surrogate models for history matching.
Figure 12: Box plots of mean absolute saturation errors with different thresholds ($\varepsilon$) for interpreted seismic surrogate (Seismic) and monitoring well surrogate (Monitor) over 500 test-case models.
5.3 Robustness of surrogates towards noise
In the previous subsection, we evaluated the performance of our surrogate models and showed that they provide a high degree of accuracy relative to the underlying simulation results. It is also useful to evaluate surrogate model robustness to ‘noise’ in the input geomodel. In such cases the geomodel will not exactly honor the assumed parameterization $\mathbf{m}=\mathbf{m}(\mathbf{h},\boldsymbol{\xi})$. For this assessment, we add Gaussian noise with zero mean and a standard deviation of 10% of the maximum value to the $\log k$, $\phi$, and $\log_{10}a_{r}$ values in all cells in a set of test-case realizations. Two sets of realizations (original, using $\mathbf{m}=\mathbf{m}(\mathbf{h},\boldsymbol{\xi})$, and perturbed) are shown in Figure 13. The quantity displayed is $\log k$ in an $x$-$z$ cross-section of the model. Fine-scale noise is clearly evident in the perturbed realizations.
Next, using the surrogate trained on realizations satisfying $\mathbf{m}=\mathbf{m}(\mathbf{h},\boldsymbol{\xi})$, we predict seismic saturation fields for both the original and perturbed models. These results are displayed in Figure 14, along with (filtered) GEOS simulation results for the original model. We see that the predictions for the perturbed models are almost identical to those for the original models, which demonstrates the robustness of the surrogate to noise (and small variations) in the input geomodel. Similar results are also observed for the monitoring well surrogate, though for brevity these are not presented here.
(a) Realization 1 (original)
(b) Realization 1 (perturbed)
(c) Realization 2 (original)
(d) Realization 2 (perturbed)
Figure 13: Original and perturbed geomodels. The quantity displayed is $\log k$ for the $x$-$z$ cross section through the injection well.
(a) Real. 1 (sim, original $\mathbf{m}$)
(b) Real. 1 (surr, original $\mathbf{m}$)
(c) Real. 1 (surr, perturbed $\mathbf{m}$)
(d) Real. 2 (sim, original $\mathbf{m}$)
(e) Real. 2 (surr, original $\mathbf{m}$)
(f) Real. 2 (surr, perturbed $\mathbf{m}$)
Figure 14: Comparison of reference simulation results (sim, original $\mathbf{m}$) and surrogate predictions with original geomodel (surr, original $\mathbf{m}$) and perturbed geomodel (surr, perturbed $\mathbf{m}$). Results are for interpreted seismic saturation fields in the $x$-$z$ cross section through the injection well at 1 year for two different realizations.
@@ 6.1 Setup, errors and likelihood function
The true (synthetic) geomodel, used to provide observed data, is a realization characterized by the following metaparameter values: $\mu_{\log k}=3.0$, $\sigma_{\log k}=2.3$, $\log_{10}a_{r}=-1.8$ ($a_{r}=0.0158$), $d=0.03$ and $e=0.055$. The true-model response for CO2 saturation, along $y$-$z$ and $x$-$z$ cross sections through the injection well at a time of 1 year, is presented in Figure 15(a) and (b). The corresponding filtered seismic interpretations are displayed in Figure 15(c) and (d). The filtered response loses some of the detail, but the heterogeneity-driven channeling features still appear. The plume locations for this case are presented in Figure 15(e) and (f). Note that the channeling is somewhat preserved, though the level of detail is clearly less than that in Figure 15(c) and (d).
@@
(e) Plume location $y$-$z$ cross section
(f) Plume location $x$-$z$ cross section
Figure 15: True-model saturation fields at 1 year. Top row shows high-fidelity results for cross sections through the injection well, second row shows corresponding filtered (interpreted saturation seismic) results, and bottom row shows plume location using a saturation threshold of 0.05.
Some amount of random measurement error (noise) will be present in the observed seismic and monitoring data. To represent this measurement error, as well as other types of error such as model resolution error, we add a noise vector $\boldsymbol{\epsilon}$ to the true data used in history matching. The observed data vector is thus expressed as:
$\mathbf{d}_{\text{obs}}=\mathbf{d}_{\text{true}}+\boldsymbol{\epsilon}.$ (13)
Here we take $\boldsymbol{\epsilon}$ to be uncorrelated and unbiased, i.e., of zero mean and diagonal covariance matrix. The measurement error and model resolution error are assigned directly in this work. For the monitoring well data, the standard deviation of the combined measurement and model resolution errors is set to be 5% of the maximum of the collected monitoring data over all available time steps. In our case this is 0.0162, which is close to the value (0.02) used by Sun and Durlofsky (2019). For the interpreted seismic data, the standard deviation of the error is set to be 10% of the maximum of the observed data over all available time steps. This is similar to the approach used by Gervais and Roggero (2010), who set the standard deviation of error to be 10% of the seismic data mean. The noise for the (synthetic) seismic data is added to the ‘true’ high-fidelity simulation results before the filtering procedure, consistent with the approach used by Bukshtynov et al. (2015). Other treatments, e.g., error varying as a function of saturation, could be readily incorporated into our framework.
Error in the surrogate models relative to full-physics simulation also enters the history matching formulation. This surrogate-model approximation error, denoted $\boldsymbol{\epsilon}_{\text{surr}}$, is estimated during the test-set evaluation. Specifically,
@@
where $\mathbf{d}_{\text{pred}}$ and $\mathbf{d}_{\text{true}}$ denote data predicted by the surrogate and true models, respectively. These errors can be computed at each time step used in the history matching. In our case, the mean error (over all time steps) in the monitoring well data surrogate is 0.00124. For the seismic surrogate, the mean error for interpreted saturation seismic and plume location seismic are -0.0001 and -0.0004, respectively. Thus we see that the average errors are near zero, consistent with the assumption of unbiased errors.
The standard deviations ($\sigma_{\text{surr}}$) of the errors are larger than the mean errors. Specifically, the observed $\sigma_{\text{surr}}$ values are around 0.03 for monitoring well data, 0.01 for interpreted seismic data, and 0.1 for plume location data. It is these quantities that enter into the history matching computations. The total standard deviation ($\sigma_{\text{total}}$), which combines the standard deviation of $\boldsymbol{\epsilon}_{\text{surr}}$ ($\sigma_{\text{surr}}$) and the measurement and model resolution errors (we denote the standard deviation of these two errors as $\sigma_{\text{d}}$), is obtained via $\sigma_{\text{total}}=\left(\sigma_{\text{surr}}^{2}+\sigma_{\text{d}}^{2}\right)^{1/2}$. In our history matching computations, the total standard deviation is calculated separately for each history matching time step (based on $\sigma_{\text{surr}}$ at that time step).
@@
We will apply the hierarchical MCMC method introduced in Section 4 for three cases – monitoring well data only, monitoring well data plus interpreted seismic plume location, and monitoring well data plus interpreted seismic saturation. This will allow us to assess the impact of different types and combinations of data on uncertainty reduction. In cases with seismic data, the likelihood appearing in Eq. 10 and Eq. 11 is given by
@@
$\begin{split}p(\mathbf{d}_{\text{obs}}\mid\boldsymbol{\xi},\mathbf{h})=c\exp\left(\frac{1}{T_{\text{mon}}}\left(-\frac{1}{2}(\mathbf{d}_{\text{obs}}^{\text{mon}}-f_{\text{mon}}(\boldsymbol{\xi},\mathbf{h}))^{T}C_{\text{mon}}^{-1}(\mathbf{d}_{\text{obs}}^{\text{mon}}-f_{\text{mon}}(\boldsymbol{\xi},\mathbf{h}))\right.\right)\\ +\frac{1}{T_{\text{seis}}}\left(\left.-\frac{1}{2}(\mathbf{d}_{\text{obs}}^{\text{seis}}-f_{\text{seis}}(\boldsymbol{\xi},\mathbf{h}))^{T}C_{\text{seis}}^{-1}(\mathbf{d}_{\text{obs}}^{\text{seis}}-f_{\text{seis}}(\boldsymbol{\xi},\mathbf{h}))\right)\right),\end{split}$ (15)
@@
where $c$ is a normalization constant, $\mathbf{d}_{\text{obs}}^{\text{mon}}$ and $\mathbf{d}_{\text{obs}}^{\text{seis}}$ denote observed monitoring well data and interpreted seismic data, respectively, and $C_{\text{mon}}$ and $C_{\text{seis}}$ are the total error covariance matrices for monitoring well and seismic data, respectively. The parameters $T_{\text{mon}}$ and $T_{\text{seis}}$, referred to as simulated temperatures, control the impact of the different data types and the MCMC acceptance rate (Li, 2012). If only monitoring data are used for history matching, the seismic data portion of the likelihood does not appear, resulting in
@@
$\begin{split}p(\mathbf{d}_{\text{obs}}\mid\boldsymbol{\xi},\mathbf{h})=c\exp\left(\frac{1}{T_{\text{mon}}}\left(-\frac{1}{2}(\mathbf{d}_{\text{obs}}^{\text{mon}}-f_{\text{mon}}(\boldsymbol{\xi},\mathbf{h}))^{T}C_{\text{mon}}^{-1}(\mathbf{d}_{\text{obs}}^{\text{mon}}-f_{\text{mon}}(\boldsymbol{\xi},\mathbf{h}))\right)\right).\end{split}$ (16)
Based on numerical experimentation, we specify $T_{\text{mon}}=0.8$ when only monitoring data are used, $T_{\text{mon}}=1.5$ and $T_{\text{seis}}=200$ when monitoring well data and plume location seismic data are used, and $T_{\text{mon}}=1$ and $T_{\text{seis}}=6$ when monitoring well data and interpreted saturation seismic data are used. These settings result in MCMC acceptance rates in the appropriate range, i.e., 10% to 40% (Gelman et al., 1996; Han et al., 2024), and balance the impact of the different terms in the likelihood function.
@@
Additional parameter specifications are as follows. We set $\beta=0.05$, where $\beta$ (appearing in Eq. 9) scales the magnitude of sample updates. The standard deviation of the Gaussian proposal distribution for the metaparameters is set to be 0.05 of the parameter range. In the results below, hierarchical MCMC using only monitoring well data requires 57,819 function evaluations (17,750 sets of metaparameters and their associated PCA latent variables are accepted). Results using monitoring well data and plume location seismic data require 83,353 function evaluations (17,250 sets of metaparameters and associated PCA variables accepted). Results using monitoring well data and saturation seismic data require 61,592 function evaluations (15,000 sets of metaparameters and associated PCA variables accepted).
Recall that, to train the surrogate models, we required 4000 simulation runs and about 2.25 hours of GPU-based training. Once these tasks are completed, the MCMC history matching for each case (involving $\sim$58,000 or more function evaluations plus associated geomodel construction computations) can be completed relatively quickly, in about 9 hours. Without these deep learning-based surrogate models, the history matching framework applied in this study would be intractable.
Prior to performing history matching, it is useful to conduct a sensitivity analysis to quantify the impact of the various parameters on the data types used in data assimilation. We performed such a sensitivity analysis, of the type presented by Han et al. (2024), for our case. The findings are that, with either monitoring well or seismic saturation data, the most sensitive parameters are the PCA variables, $\boldsymbol{\xi}$. Monitoring well data are then most sensitive to $\mu_{\log k}$, $\sigma_{\log k}$ and $a_{r}$; very little sensitivity to $d$ and $e$ is observed. Seismic data show some sensitivity to $\mu_{\log k}$ and $\sigma_{\log k}$, but they are less sensitive to the other parameters. Thus, we expect the data to be informative in terms of $\boldsymbol{\xi}$ (i.e., particular realizations) and somewhat informative for the metaparameters other than $d$ and $e$. These general expectations will be seen to be consistent with our history matching results.
@@ 6.2 History matching results
Posterior results for the metaparameters, obtained from the last 10,000 accepted samples, for cases both with and without seismic data, are presented in Figures 16 and 17. The gray region in each figure represents the prior distribution of the metaparameter and the red dashed line indicates the true value of the metaparameter. The posterior histograms appear in blue for results in which only monitoring well data are used, in green for results where monitoring well data and interpreted plume location data are used, and in orange for results where monitoring well data and interpreted seismic saturation data are used. A reasonable degree of uncertainty reduction is achieved for $\sigma_{\log k}$ and $\log_{10}a_{r}$ when only monitoring well data are used (Figure 16(d) and (g)). In addition, it is evident that the mode of the posterior histogram is near the true value for these quantities. We observe a small amount of uncertainty reduction for $\sigma_{\log k}$ (Figure 16(a)) in this case.
Very little uncertainty reduction is observed for parameters $d$ and $e$ in this case (Figure 17(a) and (d)), indicating that the monitoring well data are not informative in terms of the relationship between $\phi$ and $\log k$. This is consistent with the observations of Han et al. (2024). A small degree of uncertainty reduction is achieved, however, in the porosity field itself. This is demonstrated in Figure 17(g), where we show the prior and posterior distributions for the mean of the porosity, $\mu_{\phi}$. The results in Figure 17(g) are determined by using the relationship between $\phi$ and $\log k$ in Eq. 2.
Considerably more uncertainty reduction in the key metaparameters is achieved when both monitoring well data and seismic data are used. This is the case using either interpreted plume location or seismic saturation data. Specifically, with both monitoring well data and plume location data, we observe narrower distributions, which include the true value, for $\mu_{\log k}$, $\sigma_{\log k}$ and $\log_{10}a_{r}$ (Figure 16(b), (e), (h)). The posterior distributions for $\sigma_{\log k}$ and $\log_{10}a_{r}$ are narrower, and include the true value, when monitoring well data and interpreted saturation seismic data are used (Figure 16(f) and (i)). Interestingly, for $\mu_{\log k}$ there is more uncertainty reduction using plume location data, though the distribution is more centered around the true value for the case with seismic saturation data. In Figure 16(i), the maximum frequency is observed at the edge of the allowable range for the metaparameter. This (slightly) anomalous behavior may be due to correlated errors or to the small biases that exist in the surrogate models. In any event, the true value of $a_{r}$ is 0.0158, and the minimum allowable value is 0.01, so these values are close.
There is also more uncertainty reduction in $d$ and $e$ in the cases with seismic data (Figure 17(b), (c), (e), and (f)), though significant posterior uncertainty in these quantities remains. As is apparent in Figure 17(h), the use of both monitoring well data and plume location data provides substantial uncertainty reduction in mean porosity itself. A slightly narrower histogram is observed with interpreted saturation data (compare Figure 17(h) and (i). Taken in total, the results in Figures 16 and 17 clearly show the additional (significant) uncertainty reduction achieved when interpreted seismic data are incorporated into the data assimilation procedure, even if these data provide only plume location rather than interpreted saturation.
@@
(b) $\mu_{\log k}$ with plume location
(c) $\mu_{\log k}$ with interpreted $S_{g}$
(d) $\sigma_{\log k}$ without seismic
(e) $\sigma_{\log k}$ with plume location
(f) $\sigma_{\log k}$ with interpreted $S_{g}$
(g) $\log_{10}a_{r}$ without seismic
(h) $\log_{10}a_{r}$ with plume location
(i) $\log_{10}a_{r}$ with interpreted $S_{g}$
Figure 16: History matching results for metaparameters $\mu_{\log k}$, $\sigma_{\log k}$ and $\log_{10}a_{r}$ with monitoring well data only (left column, blue histograms), with monitoring well data and interpreted plume location seismic data (middle column, green histograms), and with monitoring well data and interpreted saturation seismic data (right column, orange histograms). Priors shown in gray and true values indicated by the vertical red line.
@@
(b) $d$ with plume location
(c) $d$ with interpreted $S_{g}$
(d) $e$ without seismic
(e) $e$ with plume location
(f) $e$ with interpreted $S_{g}$
(g) $\mu_{\phi}$ without seismic
(h) $\mu_{\phi}$ with plume location
(i) $\mu_{\phi}$ with interpreted $S_{g}$
Figure 17: History matching results for metaparameters $d$ and $e$ and average porosity $\mu_{\phi}$ with monitoring well data only (left column, blue histograms), with monitoring well data and interpreted plume location seismic data (middle column, green histograms), and with monitoring well data and interpreted saturation seismic data (right column, orange histograms). Priors shown in gray and true values indicated by the vertical red line.
We next present saturation fields (CO2 plumes) for prior and posterior geomodels. In order to identify ‘representative’ interpreted seismic saturation fields, and thus enable meaningful comparisons, we proceed as follows. We randomly select 1000 prior geomodel realizations from the training set, 1000 accepted realizations using only monitoring well data, 1000 accepted realizations using monitoring well data and plume location seismic data, and 1000 accepted realizations using monitoring well data and saturation seismic data. Interpreted seismic data are then generated for each of these 4000 realizations using the seismic surrogate model. Then, for each set of 1000 models, we apply k-means clustering to construct four clusters. The medoid from each cluster, identified with a k-medoids method, is then viewed as a representative interpreted seismic saturation field.
The resulting 16 saturation fields are shown in Figure 18 ($y$-$z$ cross sections through the injection well) and in Figure 19 ($x$-$z$ cross sections through the injection well). These results are all at a time of 1 year. The true results for this case are shown in Figure 15(c) and (d). The plumes in the prior realizations are fairly regular, varying from near-cylindrical to cone-shaped. The plumes for the posterior models all differ considerably from any of the prior plumes, and all display some amount of CO2 channeling toward the bottom of the model.
Posterior saturation fields in which seismic data (of either type) are used are closer to the true saturation fields than those obtained with just monitoring well data. For example, the plumes in Figure 18(e)-(h) are limited in extent in some of the middle layers, which is not consistent with the true result in Figure 15(c). With either type of seismic data (Figure 18(i)-(p)), the saturation fields more closely resemble the true results, both in terms of the shape of the CO2 plume in the middle layers and the channeling toward the bottom of the model. Posterior results obtained with interpreted saturation seismic data appear to provide slightly more realistic saturation fields than those using only plume location data. For example, in Figure 18(m)-(p), we see three somewhat distinct channels to the left of the injector along with CO2 moving further to the left than the right. These behaviors, consistent with Figure 15(c), are not consistently observed in Figure 18(i)-(l). Analogous observations can be made in the $x$-$z$ cross sections shown in Figure 19(m)-(p) in relation to those in Figure 19(i)-(l), i.e., the extensive channeling at the bottom of the model evident in Figure 15(d) is better resolved in Figure 19(m)-(p).
@@
(i) Post. 1 (w/ loc seis)
(j) Post. 2 (w/ loc seis)
(k) Post. 3 (w/ loc seis)
(l) Post. 4 (w/ loc seis)
(m) Post. 1 (w/ $S_{g}$ seis)
(n) Post. 2 (w/ $S_{g}$ seis)
(o) Post. 3 (w/ $S_{g}$ seis)
(p) Post. 4 (w/ $S_{g}$ seis)
Figure 18: Representative interpreted seismic saturation fields in $y$-$z$ cross sections through the injector at 1 year for prior realizations (first row), posterior realizations obtained using only monitoring well data (second row), posterior realizations obtained using monitoring well and plume location seismic data (third row), and posterior realizations obtained using monitoring well and saturation seismic data (fourth row). True result shown in Figure 15(c).
@@
(i) Post. 1 (w/ loc seis)
(j) Post. 2 (w/ loc seis)
(k) Post. 3 (w/ loc seis)
(l) Post. 4 (w/ loc seis)
(m) Post. 1 (w/ $S_{g}$ seis)
(n) Post. 2 (w/ $S_{g}$ seis)
(o) Post. 3 (w/ $S_{g}$ seis)
(p) Post. 4 (w/ $S_{g}$ seis)
Figure 19: Representative interpreted seismic saturation fields in $x$-$z$ cross sections through the injector at 1 year for prior realizations (first row), posterior realizations obtained using only monitoring well data (second row), posterior realizations obtained using monitoring well and plume location seismic data (third row), and posterior realizations obtained using monitoring well and saturation seismic data (fourth row). True result shown in Figure 15(d).